### v1

# Wilsonian Renormalization of Neural Network Gaussian Processes

Abstract
Separating relevant and irrelevant information is key to any modeling process or scientific inquiry. Theoretical physics offers a powerful tool for achieving this in the form of the renormalization group (RG). Here we demonstrate a practical approach to performing Wilsonian RG in the context of Gaussian Process (GP) Regression. We systematically integrate out the unlearnable modes of the GP kernel, thereby obtaining an RG flow of the Gaussian Process in which the data plays the role of the energy scale. In simple cases, this results in a universal flow of the ridge parameter, which becomes input-dependent in the richer scenario in which non-Gaussianities are included. In addition to being analytically tractable, this approach goes beyond structural analogies between RG and neural networks by providing a natural connection between RG flow and learnable vs. unlearnable modes. Studying such flows may improve our understanding of feature learning in deep neural networks, and identify potential universality classes in these models.
1 Introduction
The renormalization group (RG) is one of the pillars of theoretical physics peskin1995 ; cardy1996scaling . Besides providing an analytical toolkit for tackling large-scale interacting systems (e.g. WILSON1974 ; Gross1973 ; KT1973 ), RG underpins concepts such as universality, irrelevance, and scaling. In complex physical systems, RG is crucial for our ability to model and predict emergent phenomena. It is therefore interesting to ask how may this conceptual framework generalizes to complex virtual systems, in particular deep neural networks (DNNs).
From an experimental point of view, DNNs exhibit various phenomena common to RG domains. They are large-scale systems kaplan2020scaling , exhibiting collective phenomena where the importance of each specific parameter is minute molchanov2017pruning . In realistic settings, DNNs often show robust scaling laws in performance versus dataset size or the number of tunable parameters bahri2021explaining ; kaplan2020scaling . In a related manner, data-data covariance matrices, associated with networks’ outputs, also exhibit power law decaying spectra bahri2021explaining ; bordelon2021spectrum similar in spirit to Zipf’s law. Finally, universality seems present to some degree, as often large microscopic/mechanistic changes, such as a change of activation function or loss function, result in minor performance changes Novak2018BayesianDC ; Lee2020FiniteVI .
Here we suggest a distinct route where RG may come into quantitative use. To this end, we model DNNs as Gaussian Processes (GPs), a description which becomes exact in a suitable infinite-overparametrization limit Neal ; jacot2018neural ; lee2018deep ; matthews2018gaussian ; Cohen_2021 and applies to a wide range of DNNs. It is also extendable to finite DNNs via adaptive kernel approaches LiSompolinsky2021 ; Seroussi2023 ; ariosto2022statistical . In the GP viewpoint, the action governing the theory is quadratic and described by a kernel which encodes the network’s architecture. Diagonalizing this kernel, one obtains a set of eigenfunctions. Fluctuations of network outputs along eigenfunctions with low kernel eigenvalues are highly suppressed, and are thus analogous to “high-energy” modes in RG. [^1] To avoid confusion, note that the language “high” and “low energy” stems from the field-theoretic practice of working in frequency space; here, the kernel eigenmodes are akin to wavelengths, i.e., “high-energy” modes mean modes with low eigenvalues. Such modes, if present in the target, are also harder to learn Rasmussen ; Cohen_2021 ; canatar . Notably, while the GP theory is quadratic and, in this sense, exactly solvable, it contains randomness (or quenched disorder) associated with the specific choice of training set. Averaging over the latter leads to a non-quadratic/interacting replicated field theory Cohen_2021 ; canatar .
Our contributions.
In this work, we apply Wilsonian RG to Gaussian Processes regression by iteratively integrating out these high-energy GP modes. Sec. 2 presents general layout for this approach. We show that in various relevant scenarios, such a procedure is tractable and results in a renormalized GP theory. When the high-energy eigenfunctions are uncorrelated, in a manner to be made precise below, this amounts to a scalar renormalization of the loss or equivalently the ridge parameter; this is developed in sec. 3. When eigenfunction correlations are present, a further input-dependent/spatial weighting of the loss appears as well as potential shifts to the target; this is demonstrated in sec. 4. The resulting RG flow proceeds until it reaches the first learnable eigenfunction. Thus, our lower RG scale (i.e., the infrared) naturally depends on the type and amount of data. In cases where high-energy modes have sizable collective fluctuations, the RG flow ends in a weakly interacting regime amenable to perturbative methods Cohen_2021 . Alternatively, it may end in a strongly interacting theory which, in generic high dimensional scenarios, is amenable to saddle point techniques canatar . Lastly, we point out that for power-law decaying kernel spectra, often found in real-world data, we obtain an RG flow towards the weakly interacting regime. By relying on rather general formalism and introducing a precise and practically motivated notion of RG, we hope to provide a stepping stone towards establishing notions of universality in deep learning.
Related works.
Several works have considered connections and analogies between deep learning and RG. For example, Halverson:2020trp ; Erbin_2022 ; erbin2022renormalization ; erbin2023functional examined RG flows in the couplings of a corresponding field theory due to finite-width effects (see also Grosvenor:2021eol ). Similar effects arising layer-by-layer have been discussed in Roberts_2022 from an effective theory perspective. The structural analogy between RG and DNNs was further explored in Erdmenger:2021sot , which tracked the flow of information as a function of depth. These approaches are however primarily concerned with untrained networks and do not provide a relation between RG and learning dynamics via the data. Another line of work Cotler_2023 ; cotler2023renormalizing explored RG-inspired training schemes for diffusion models and optimal transport in DNNs, where updates of network outputs are modeled after Polchinski’s so-called Exact RG (ERG) flow. In addition, berman2022dynamics ; berman2022inverse ; Berman_2023 ; berman2024ncoder construct an information theoretic ERG flow to describe the inverse of Bayesian inference in DNN ensembles. What distinguishes our approach is the natural connection between the energy scale in Wilsonian RG and the data, which allows us to perform RG on the network’s degrees of freedom for networks trained in a standard manner, using the framework in Cohen_2021 .
Finally, as this manuscript was completed, a related work atanasov2024scaling appeared where an approach based on free probability was used to derive “renormalization” of the ridge parameter and other quantities relevant for GP regression with Gaussian features. There, the change to the ridge parameter arises from a self-energy-type diagrammatic resummation; however, while diagrammatic resummations are of course part of the toolkit of field theory, central RG notions such as a cut-off or a flow of couplings were not made explicit. The current manuscript goes beyond the case of Gaussian features, and develops a direct Wilsonian RG framework which further generalizes to finite-width effects systematically.
2 Equivalent Kernel Limit of the Replica Action
To provide a relatively self-contained derivation, we first define Gaussian Process Regression, its relations to neural networks, and replica techniques used to obtain the average predictions over ensembles of datasets. In addition to being a well-established tool in machine learning, Gaussian Process Regression describes both the lazy learning or kernel regime of an infinite-width neural network lee2018deep ; jacot2018neural ; Cohen_2021 , alongside extensions to finite networks via the aforementioned adaptive kernel approaches.
The output of such random networks behaves as a Gaussian Process with zero mean and a covariance function $K(x,x^{\prime})$, called the kernel. Here, the training data, $x,x^{\prime}\in\mathbb{R}^{d}$, are drawn from a data measure [^2] By this we simply mean a measure over the data that admits an interpretation as a probability distribution. That is, let $\mathrm{d}\mu_{x}$ be a measure on $x\in\mathbb{R}^{d}$ such that for some probability distribution $p_{\mathrm{data}}(x)$, we have $\int\!f(x)\mathrm{d}\mu_{x}=\int\!f(x)p_{\mathrm{data}}(x)\mathrm{d}x$ for $f:\mathbb{R}^{d}\to\mathbb{R}$. In a slight abuse of notations, we will refer to both the measure $\mathrm{d}\mu_{x}$ and the distribution $p_{\mathrm{data}}(x)$ as the data measure. $p_{\mathrm{data}}(x)$; the total training set consists of $n$ such draws, and is denoted by the $n\times d$ matrix $X_{n}$. We further consider a regression problem with a target, $y(x)$. A standard result in GP regression states that Rasmussen the average predictor on a test point, $\tilde{f}(x_{*})$, given $X_{n}$ is
$\tilde{f}(x_{*}|X_{n})=K(x_{*},X_{n})[K(X_{n},X_{n})+\sigma^{2}\mathbb{1}_{n\times n}]^{-1}y(X_{n})\leavevmode\nobreak\ ,$ (1)
where $K(X_{n},X_{n})$ is understood as an $n\times n$ matrix, and $\sigma^{2}$ is the variance of the observation noise, called the ridge parameter. (That is, we wish to model more realistic situations in which we do not have access to the function values $y(x)$ themselves, but only to noisy versions thereof, $y(x)+\epsilon$, where $\epsilon$ is additive Gaussian noise with variance $\sigma^{2}$.) The latter sets an overall tolerance to discrepancies between the target and the prediction on the training set. As we shall see, one main effect of RG is to renormalize this tolerance and also, in more advanced settings (see below), make the ridge parameter input-dependent.
Since the average predictor (1) is conditional on $X_{n}\!\sim\!p_{\mathrm{data}}(x)$, it is sensitive to the details of the specific dataset we drew. To remove this dependence, we consider the dataset-averaged predictor by computing the expectation value with respect to $p_{\mathrm{data}}(x)$. The most general way of doing so in a tractable manner uses a “grand-canonical” trick Malzahn ; Cohen_2021 , which assumes that the number of training points is drawn from a Poisson distribution with expectation value $\eta$,
$\bar{f}(x_{*})=\sum_{n=1}^{\infty}\frac{\eta^{n}e^{-\eta}}{n!}\langle\tilde{f}(x_{*}|X_{n})\rangle_{X_{n}\sim p_{\mathrm{data}}(x)}\eqqcolon\langle\tilde{f}(x_{*}|X_{n})\rangle_{\eta}\leavevmode\nobreak\ ,$ (2)
such that $\eta$ now controls the average number of data points. Notably, due to the nature of the Poisson distribution, in the relevant limit of large $\eta$, the fluctuations become increasingly negligible compared to the average.
In statistical physics, (2) is known as a quenched average. Such averages are generally written in terms of the free energy as $\partial_{\alpha}\langle\ln(Z_{\alpha})\rangle_{X\sim p_{\mathrm{data}}(x)}\big{|}_{\alpha=0}$, where $Z$ is the partition function and $\alpha$ is some source field that couples to $f(x)$. [^3] In field-theoretic language, this is just the connected correlation function generated by $\ln Z$. We will suppress the source fields $\alpha$, as they play no role in the rest of the setup. However, the presence of the logarithm makes explicitly computing this challenging in all but the simplest cases. To this end, a common solution is to use the replica trick, in which we take $M$ copies of the system, and recover $\ln Z$ in the limit $M\to 0$ via
$\ln Z=\lim_{M\to 0}\frac{1}{M}\!\left(Z^{M}-1\right)\leavevmode\nobreak\ .$ (3)
We can now perform the quenched average on the $M$-copy system $Z^{M}$ instead, which is often more tractable.
This replica approach was first applied to GP regression in Malzahn , and was used in Cohen_2021 to obtain
$\langle Z^{M}\rangle_{\eta}=e^{-\eta}\int\!\prod_{m=1}^{M}\mathcal{D}f_{m}e^{-S}\leavevmode\nobreak\ ,$ (4)
where the action is
$S=\sum_{m=1}^{M}\frac{1}{2}\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\,f_{m}(x)K^{-1}(x,x^{\prime})f_{m}(x^{\prime})-\eta\int\!\mathrm{d}\mu_{x}\,e^{-\sum_{m=1}^{M}\frac{\left(f_{m}(x)-y(x)\right)^{2}}{2\sigma^{2}}}\leavevmode\nobreak\ ,$ (5)
[…]
Although our initial field theory was Gaussian, averaging over the datasets resulted in a non-quadratic action, due to “quenched-disorder" induced by the training set. A simple yet useful limit Sollich2004 of the above theory, known as the Equivalent Kernel limit, is obtained when $\eta,\sigma^{2}\rightarrow\infty$ with $\eta/\sigma^{2}$ fixed. In this case the exponential on the right-hand side of (5) can be Taylor expanded, so that, at leading order, we obtain a non-interacting theory,
$\eta+S=\sum_{m=1}^{M}\frac{1}{2}\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\,f_{m}(x)K^{-1}(x,x^{\prime})f_{m}(x^{\prime})+\frac{\eta}{\sigma^{2}}\int\!\mathrm{d}\mu_{x}\sum_{m}\frac{\left(f_{m}(x)-y(x)\right)^{2}}{2}\leavevmode\nobreak\ .$ (6)
It is then convenient to write this in terms of the eigenfunctions, $\phi_{k}(x)$, of the kernel $K$ with respect to the data measure $\mathrm{d}\mu_{x}=p_{\mathrm{data}}(x)dx$. That is, the aforementioned covariance function admits an eigensystem decomposition
$\int\!\mathrm{d}\mu_{x}\,K(x,x^{\prime})\phi_{k}(x)=\lambda_{k}\phi_{k}(x^{\prime})\leavevmode\nobreak\ ,$ (7)
where $\phi_{k}$ is the kernel eigenfunction with eigenvalue $\lambda_{k}$ which form a complete orthonormal basis, i.e.,
$\int\!\mathrm{d}\mu_{x}\,\phi_{k}(x)\phi_{q}(x)=\delta_{kq}\leavevmode\nobreak\ .$ (8)
This allows us to make the following spectral decomposition (analogous to Fourier transforming to momentum space),
$f_{m}(x)=\sum^{\infty}_{k=1}f_{mk}\phi_{k}(x)\leavevmode\nobreak\ ,\qquad\quad y(x)=\sum^{\infty}_{k=1}y_{k}\phi_{k}(x)\leavevmode\nobreak\ .$ (9)
We refer to the coefficients $f_{mk}$ as the GP modes, and the functions $\phi_{k}$ as the feature modes. Note that the former carry replica indices and may be highly correlated, while the feature modes are an inert property of the kernel/lazy-network. Inserting these decompositions into the quadratic action (6) yields the decoupled “feature-space” action
$S=\frac{1}{2}\sum_{m=1}^{M}\sum_{k=1}^{\infty}\left(\frac{1}{\lambda_{k}}|f_{mk}|^{2}+\frac{\eta}{\sigma^{2}}|f_{mk}-y_{k}|^{2}\right)\leavevmode\nobreak\ ,$ (10)
see Appendix A. Note that the path integral measure also decouples into a measure over the feature modes, $\mathcal{D}f_{m}\sim\prod_{k}\mathcal{D}f_{mk}$, up to a constant Jacobian factor. Given this action, one can then show that the average prediction is Cohen_2021 ; Rasmussen
$\bar{f}_{k}=\frac{\lambda_{k}}{\lambda_{k}+\sigma^{2}/\eta}\,y_{k}\leavevmode\nobreak\ ,$ (11)
with dataset-averaged GP fluctuations given by
$\mathrm{Var}[f_{k}]=\frac{1}{\lambda_{k}^{-1}+\eta\,\sigma^{-2}}=\frac{\sigma^{2}}{\eta}\frac{\lambda_{k}}{\lambda_{k}+\sigma^{2}/\eta}\leavevmode\nobreak\ .$ (12)
In the context of our RG prescription introduced below, these expression have the following important interpretation: modes with $\lambda_{k}\ll\sigma^{2}/\eta$ essentially behave as if $\sigma^{2}=\infty$ or $\eta=0$, i.e., they are decoupled from the inference problem and hence could just as well have been integrated out. This fact is already implicit in the action (10), where the term proportional to $\eta/\sigma^{2}$ becomes negligible compared to $1/\lambda_{k}$ for sufficiently small $\lambda_{k}$.
3 Renormalizing Gaussian features
Formally, our Wilsonian renormalization procedure can be explained very simply: we wish to integrate out all modes above some cutoff $\kappa$, in order to obtain an effective action in terms of only the modes with $k<\kappa$. Accordingly, we split the spectral decomposition (9) into
$f_{m<}(x)\coloneqq\sum_{k\leq\kappa}f_{mk}\phi_{k}(x)\leavevmode\nobreak\ ,\qquad\quad f_{m>}(x)\coloneqq\sum_{k>\kappa}f_{mk}\phi_{k}(x)\leavevmode\nobreak\ ,$ (13)
and similarly
$y_{<}(x)\coloneqq\sum_{k\leq\kappa}y_{k}\phi_{k}(x)\leavevmode\nobreak\ ,\qquad\quad y_{>}(x)\coloneqq\sum_{k>\kappa}y_{k}\phi_{k}(x)\leavevmode\nobreak\ .$ (14)
We refer to the coefficients $f_{mk}$ as the GP modes, and the functions $\phi_{k}$ as the feature modes. Note that the former carry replica indices and may be highly correlated, while the feature modes are an inert property of the kernel/lazy-network. In the Equivalent Kernel limit, one sees that the action (10) splits cleanly into greater and lesser modes, $S=S_{0}[f_{m<}]+S_{0}[f_{m>}]$. At finite $\eta,\sigma^{2}$ however, the replicated action (5) will contain interaction terms collectively denoted $S_{\mathrm{int}}[f_{m<},f_{m>}]$. Explicitly, making the decomposition (13), (14) in the replicated action (5), we have
$S=$ $\frac{1}{2}\sum_{m=1}^{M}\left(\sum_{k=1}^{\kappa}\frac{1}{\lambda_{k}}|f_{mk}|^{2}+\sum_{k=\kappa+1}^{\infty}\frac{1}{\lambda_{k}}|f_{mk}|^{2}\right)$ (15)
$-\eta\int\!\mathrm{d}\mu_{x}\,\exp\!\left(-\frac{1}{2\sigma^{2}}\sum_{m=1}^{M}\left[\left(f_{m<}-y_{<}\right)+\left(f_{m>}-y_{>}\right)\right]^{2}\right)\leavevmode\nobreak\ ,$
where the first line and second line respectively define $S_{\mathrm{0}}[f_{m<}],\leavevmode\nobreak\ S_{\mathrm{0}}[f_{m>}],$ and $S_{\mathrm{int}}$. In general then, the exact partition function (4) takes the form
$\langle Z^{M}\rangle_{\eta}$ $=e^{-\eta}\int\!\prod_{m}\mathcal{D}f_{m<}\,e^{-S_{0}[f_{m<}]}\int\prod_{m}\mathcal{D}f_{m>}\,e^{-S_{0}[f_{m>}]-S_{\mathrm{int}}[f_{m<},f_{m>}]}$ (16)
$=e^{-\eta}\int\prod_{m}\mathcal{D}f_{m<}\,e^{-S_{\mathrm{eff}}[f_{m<}]}\leavevmode\nobreak\ ,$
where we have defined the effective action
$S_{\mathrm{eff}}[f_{m<}]\coloneqq-\ln\left(e^{-S_{0}[f_{m<}]}\int\prod_{m}\mathcal{D}f_{m>}\,e^{-S_{0}[f_{m>}]-S_{\mathrm{int}}[f_{m<},f_{m>}]}\right)\leavevmode\nobreak\ ,$ (17)
with $\mathcal{D}f_{m<}\coloneqq\prod_{k\leq\kappa}\mathrm{d}f_{mk}$, and similarly for $\mathcal{D}f_{m>}$. In field-theoretical language, the effective action corresponds to the low-energy (cf. footnote 1) effective action that describes the system up to the scale set by the cutoff, $\kappa$. From a Bayesian perspective, this amounts to marginalizing over the hidden degrees of freedom that are inaccessible to the observer.
Our objective is now to perform the integral over the higher modes in (17). However, the presence of the interaction term complicates this, so we will first recast this term in a way which makes the interaction between the lesser and greater GP modes more tractable. To this end, we observe that the feature modes, $\phi_{k}(x)$, are functions over the data, which in turn is drawn from the data measure, $\mathrm{d}\mu_{x}=p_{\mathrm{data}}(x)\mathrm{d}x$. It will prove more convenient to instead treat these as fluctuating random variables over the data measure. To do this, we insert an identity of the given form into the integral over the data measure $\mathrm{d}\mu_{x}$ in the interaction term
$1=\int\!\prod_{k=1}^{\infty}\mathrm{d}\varphi_{k}\,\delta\left(\varphi_{k}-\phi_{k}(x)\right)\leavevmode\nobreak\ ,$ (18)
where $\varphi_{k}$ is a random variable over data space. The interaction term then becomes
$S_{\mathrm{int}}$ $=-\eta\int\!\prod_{k=1}^{\infty}\mathrm{d}\varphi_{k}\underbrace{\int\!\mathrm{d}\mu_{x}\,\delta\left(\varphi_{k}-\phi_{k}(x)\right)}_{P[\varphi_{k}]}\exp\!\left\{-\frac{1}{2\sigma^{2}}\sum_{m=1}^{M}\left[\sum_{k=1}^{\infty}\left(f_{mk}-y_{k}\right)\phi_{k}(x)\right]^{2}\right\}$ (19)
$=-\eta\int\!\prod_{k=1}^{\infty}\mathrm{d}\varphi_{k}\,P[\varphi_{k}]\,\exp\!\left\{-\frac{1}{2\sigma^{2}}\sum_{m=1}^{M}\left[\sum_{k=1}^{\infty}\left(f_{mk}-y_{k}\right)\varphi_{k}\right]^{2}\right\}\leavevmode\nobreak\ ,$
where we have introduced the feature space distribution
$P[\varphi_{k}]\coloneqq\int\!\mathrm{d}\mu_{x}\,\delta\left(\varphi_{k}-\phi_{k}(x)\right)=\int\!\mathrm{d}x\,p_{\mathrm{data}}(x)\delta\left(\varphi_{k}-\phi_{k}(x)\right)\leavevmode\nobreak\ .$ (20)
This removes the explicit dependence on the data, so that the new feature modes $\varphi_{k}$ (not to be confused with $\phi_{k}(x)$) are considered as random variables over feature space. Note that the second line of (19) is now written entirely in terms of these variables. Intuitively, (20) integrates over the data measure $\mathrm{d}\mu_{x}$ in such a way as to ensure that the modes $\varphi_{k}$ correctly preserve the original distribution over feature space. This will enable us to capitalize on cases where $P[\varphi]$ is (approximately) Gaussian.
Analogous to (13), we then define the following two variables
$\Phi_{m<}\coloneqq\sum_{k\leq\kappa}\left(f_{mk}-y_{k}\right)\varphi_{k}\leavevmode\nobreak\ ,\qquad\quad\Phi_{m>}\coloneqq\sum_{k>\kappa}\left(f_{mk}-y_{k}\right)\varphi_{k}\leavevmode\nobreak\ .$ (21)
For a fixed set of GP modes (i.e., within the integrand of the partition function) these can be understood as random variables over feature space. Accordingly, unlike (13), these do not depend explicitly on the data, $x$. Rather than define separate variables as in (14), we have also defined these to include the targets, $y_{k}$, since these appear together, cf. above. It will also be notationally convenient to collect these into $M$-dimensional random vectors $\Phi_{>}$, $\Phi_{<}$, so that the $m^{\mathrm{th}}$ component of $\Phi_{>}$ is $\Phi_{m>}$, i.e., $[\Phi_{>}]_{m}=\Phi_{m>}$; similarly for $\Phi_{<}$. Substituting these into (19) and collecting terms, we have
$S_{\mathrm{int}}$ $=-\eta\int\!\mathcal{D}\varphi\,P[\varphi]\,\exp\!\left\{-\frac{1}{2\sigma^{2}}\sum_{m=1}^{M}\left(\Phi_{m<}+\Phi_{m>}\right)^{2}\right\}$ (22)
$=-\eta\int\!\mathcal{D}\varphi\,P[\varphi_{<},\varphi_{>}]\,\exp\!\left\{-\frac{1}{2\sigma^{2}}\left(\Phi_{<}^{\top}\Phi_{<}+\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)\right\}\leavevmode\nobreak$
$=-\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\exp\!\left\{-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right\}\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]\exp\!\left\{-\frac{1}{2\sigma^{2}}\left(\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)\right\}\leavevmode\nobreak\ ,$
where $\mathcal{D}\varphi_{<}\coloneqq\prod_{k\leq\kappa}\mathrm{d}\varphi_{k}$, and similarly for $\mathcal{D}\varphi_{>}$ (note that this is merely a shorthand notation, and does not indicate a path integral).
We can now proceed via a two-stage process, wherein we first integrate out the greater feature modes with the GP modes held fixed, and then integrate out the greater GP modes ($f_{mk}$ with $k>\kappa$). To this end, we approximate $P[\varphi]$ as a multivariate Gaussian distribution. We note in passing that a weaker requirement, namely $\Phi_{>}$ is Gaussian under the $p_{\mathrm{data}}(x)$ data measure, also suffices; cf. canatar . At a technical level, this assumption allows us to perform the spatial integration in $S_{\mathrm{int}}$ and recast the action in terms of $f_{mk}$ alone.
Recalling that the greater and lesser feature modes are by definition linearly uncorrelated under the data measure, the multivariate Gaussian approximation implies that they are also statistically independent. Consequently, $P[\varphi_{>}|\varphi_{<}]=P[\varphi_{>}]={\cal N}[0,\mathbb{1};\varphi_{>}]$. [^5] For brevity, here we assumed that $\int\mathrm{d}\mu_{x}\phi_{k>\kappa}(x)=0$, which holds in practice for many kernels. Due to the orthogonality of feature modes (cf. (8)), this also holds whenever the constant function is spanned by the lesser feature modes ($\phi_{k\leq\kappa}(x)$). The above integral over $\varphi_{>}$ now becomes Gaussian, and may be readily evaluated by expressing it in the form
$\int$ $\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]e^{-\frac{1}{2\sigma^{2}}\left(\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)}=\frac{1}{(2\pi)^{N_{>}/2}}\int\!\mathcal{D}\varphi_{>}e^{-\sum_{k>\kappa}\frac{\varphi^{2}_{k}}{2}}e^{-I_{0}[\Phi_{>}]-I_{\mathrm{int}}[\Phi_{<},\Phi_{>}]}$ (23)
$=e^{-\mathrm{tr}\ln\frac{1}{2}\left(\mathbb{1}_{N_{<}\times N_{<}}+F^{\top}F/\sigma^{2}\right)}\exp\left\{\frac{1}{2\sigma^{4}}\Phi_{<}^{\top}F\left(\mathbb{1}_{N_{<}\times N_{<}}+F^{\top}F/\sigma^{2}\right)^{-1}F^{\top}\Phi_{<}\right\}\leavevmode\nobreak\ ,$
where $F^{\top}F$ is an $N_{>}\!\times\!N_{>}$ matrix with internal (replica) indices contracted, defined via
$I_{0}[\Phi_{>}]+I_{\mathrm{int}}[\Phi_{<},\Phi_{>}]\coloneqq$ $\,\frac{1}{2\sigma^{2}}\sum_{m;kk^{\prime}>\kappa}(f_{mk}-y_{k})(f_{mk^{\prime}}-y_{k}^{\prime})\varphi_{k}\varphi_{k^{\prime}}$ (24)
$\,+\frac{1}{\sigma^{2}}\sum_{m;k\leq\kappa,k^{\prime}>\kappa}(f_{mk}-y_{k})(f_{mk^{\prime}}-y_{k}^{\prime})\varphi_{k}\varphi_{k^{\prime}}$
$\eqqcolon$ $\,\frac{1}{2\sigma^{2}}\left(\varphi_{>}^{\top}F^{\top}F\varphi_{>}+2\Phi_{<}^{\top}F\varphi_{>}\right)\leavevmode\nobreak\ .$
We can then express this result in terms of a covariance matrix of replica indices by Taylor expanding the logarithm in the coefficient and using the cyclicity of the trace to obtain
$\mathrm{tr}$ $\ln\left(\mathbb{1}_{N_{>}\times N_{>}}+F^{\top}F/\sigma^{2}\right)^{1/2}=\mathrm{tr}\left(\frac{FF^{T}}{2\sigma^{2}}-\frac{(FF^{\top})^{2}}{4\sigma^{2}}+\ldots\right)$ (25)
$=\mathrm{tr}\ln\left(\mathbb{1}_{M\times M}+FF^{\top}/\sigma^{2}\right)^{1/2}=\mathrm{tr}\ln\left(\mathbb{1}_{M\times M}+C/\sigma^{2}\right)^{1/2}\leavevmode\nobreak\ ,$
where $C\coloneqq FF^{\top}$ is an $M\!\times\!M$ covariance matrix with elements
$[C]_{mn}=(f_{mk}-y_{k})(f_{nk}-y_{k})\leavevmode\nobreak\ .$ (26)
The argument of the exponential can similarly be written in terms of this matrix by applying the Woodbury matrix identity, whereupon
$F\left(\mathbb{1}_{N_{<}\times N_{<}}+F^{\top}F/\sigma^{2}\right)^{-1}\!F^{\top}=\left(C^{-1}+\sigma^{-2}\right)^{-1}\leavevmode\nobreak\ .$ (27)
Putting everything together, the full interaction term (22) thus becomes
$S_{\mathrm{int}}$ $=-\eta\int\!\mathcal{D}\Phi_{<}P[\Phi_{<}]e^{-\frac{1}{2}\mathrm{tr}\ln\left(C/\sigma^{2}+\mathbb{1}\right)}\exp\left\{\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\left[\frac{1}{\sigma^{2}}\left(\sigma^{-2}+C^{-1}\right)^{-1}-\mathbb{1}\right]\Phi_{<}\right\}$ (28)
$=-\eta\,\,e^{-\frac{1}{2}\mathrm{tr}\ln\left(C/\sigma^{2}+\mathbb{1}\right)}\!\int\!\mathrm{d}\mu_{x}\,e^{-\frac{1}{2}\left(\bm{f}_{m<}(x)-\bm{y}_{<}(x)\right)^{\top}\left(\sigma^{2}+C\right)^{-1}\left(\bm{f}_{m<}(x)-\bm{y}_{<}(x)\right)}\leavevmode\nobreak\ ,$
where on the second line we have reversed our measure changes above, [^6] Explicitly, $\int\!\mathcal{D}\Phi_{<}P[\Phi_{<}]=\int\!\prod_{k\leq\kappa}\mathrm{d}\varphi_{k}\,P[\varphi_{k}]=\int\!\prod_{k\leq\kappa}\mathrm{d}\varphi_{k}\int\!\mathrm{d}\mu_{x}\,\delta\left(\varphi_{k}-\phi_{k}(x)\right)=\int\!\mathrm{d}\mu_{x}$. and $\bm{f}_{m<}(x)$ and $\bm{y}_{<}(x)$ are $M$-dimensional vectors with elements $f_{m<}(x)$ and $y_{<}(x)$, respectively, cf. (13) and (14). Notably, in this expression all the higher feature modes $\phi_{k>\kappa}$ have been removed, and all higher GP modes $f_{m>}$ enter only through $C$, cf. (26). We note in passing that replacing $C$ with its self-consistent mean-value leads to the same average predictor as in canatar .
Next, we recall that, as the replica limit ($M\to 0$) is taken, our partition function should provide us with the average and variance of $f_{mk}$ as obtained by GP regression on fixed data, averaged over all datasets. Notably, the average and variance are respectively first and zeroth order in $y$ per dataset and hence must remain so after this averaging. However, in contrast to this expected result, the interaction term in our setting involves arbitrarily high powers of $y$. We argue in appendix B that the expected linear behavior in $y$ is restored in the interaction term in the replica limit. An easy way to see this is would be to start with the interaction term having a scaled $\epsilon\,y_{>}$ for an infinitesimal parameter $\epsilon$, followed by divisions by the same. Any term containing second or higher order expressions in $\epsilon\,y_{>}$ would vanish, leaving aside linear terms in $y_{>}$. Following these arguments, when calculating the average predictor (which is linear in $y_{<}$), we may set $y_{>}=0$ from the start, which we do henceforth. One can then show that the contribution $\mathrm{tr}\ln\left(C/\sigma^{2}+\mathbb{1}_{M\times M}\right)$ vanishes in the replica limit. We therefore proceed with the following interaction term:
$S_{\mathrm{int}}=-\eta\!\int\!\mathrm{d}\mu_{x}\,e^{-\frac{1}{2}\left(\bm{f}_{m<}(x)-\bm{y}_{<}(x)\right)^{\top}\left(\sigma^{2}+C_{y_{>}=0}\right)^{-1}\left(\bm{f}_{m<}(x)-\bm{y}_{<}(x)\right)}\leavevmode\nobreak\ .$ (29)
We now wish to integrate out the higher GP modes still present in the covariance matrix. This is a complicated step due to nonlinear correlations between higher and lesser GP modes in (28). To proceed, we perform the Wilsonian RG procedure perturbatively as follows: we integrate out an infinitesimal momentum shell $\kappa-\delta\kappa\leq k<\kappa$, with the cutoff $\kappa$ sufficiently large such that it renders the sum of marginalized GP kernel eigenvalues $\lambda_{k}$ (cf. (7)) much smaller than the variance scale of the observation noise $\sigma^{2}$. That is, under the free theory for the higher modes, the expectation value of the GP covariance matrix (26) is [^7] To see this, observe from the first line of (15) that $\lambda_{k}$ is the variance of the Gaussian higher modes in $S_{0>}=\tfrac{1}{2}\sum_{k>\kappa}\lambda_{k}^{-1}f_{mk}^{2}$,viewed as a probability distribution in replica space. Since $[C_{y_{>}=0}]_{nm}=\sum_{k>\kappa}f_{mk}f_{nk}$, the expectation value $C_{y_{>}=0}\right>_{S_{0}}$ is the second cumulant of this distribution.
$C_{y_{>}=0}\right>_{S_{0>}}=\mathbb{1}_{M\times M}\sum_{k>\kappa}\lambda_{k}\leavevmode\nobreak\ ,$ (30)
where $\mathbb{1}_{nm}=\delta_{nm}$. We then set the cutoff such that
$\sum_{k=\kappa-\delta\kappa}^{\kappa}\lambda_{k}\eqqcolon\delta c\ll\sigma^{2}\leavevmode\nobreak\ .$ (31)
This ensures that at each step of the RG, $[C_{y_{>}=0}]_{nm}=\sum_{k}f_{mk}f_{nk}$ has a typical scale of $\delta c$. Since $\delta c/\sigma^{2}\ll 1$, we may expand the interaction term, keeping only the leading-order contributions in $C_{y_{>}=0}$. This amounts to replacing $C_{y_{>}=0}$ in (29) by its average according to the free theory, namely $C_{y_{>}=0}\rightarrow\delta c\,\mathbb{1}_{M\times M}$. We thus find a renormalized ridge parameter $\sigma^{\prime}$, given by
${\sigma^{\prime}}^{2}=\sigma^{2}+\delta c\leavevmode\nobreak\ .$ (32)
We may now repeat this procedure iteratively, as long as $\delta c$ is smaller than the effective ridge parameter. Denoting by $c$ the sum over all $\delta c$’s integrated on all shells and $\sigma^{2}_{c}$ the renormalized ridge parameter at the variance scale defined by $c$, we obtain a very simple RG flow
$\sigma_{c}^{2}=\sigma^{2}+c\leavevmode\nobreak\ .$ (33)
In the language of RG, $\sigma^{2}$ is the bare ridge parameter at which the first infinitesimal shell was defined, and $\sigma_{c}^{2}$ is the renormalized value. Of course, this procedure breaks down when the smallest $\lambda$ which remains become comparable to $\sigma_{c}^{2}$. However, long before this point, we will reach the limit $\lambda=\sigma_{c}^{2}/\eta$, which marks the threshold for learnability (cf. the discussion at the end of Sec. 2). Thus, for a generic target function, a natural point to stop the momentum shell RG flow would be at the first learnable mode.
It is interesting to compare this result with other works which use the feature Gaussianity assumption, such as canatar . There it was found that the dataset average predictor coincides with the Equivalent Kernel approximation with an effective ridge parameter, $\kappa$. This is similar to our result in that their $\kappa$ is the original ridge parameter plus all the eigenfunctions which are unlearnable. There are two important conceptual differences however: in that derivation, there is no notion of RG, nor of a Wilsonian cut-off or RG flows. The same largely holds for their concurrent work atanasov2024scaling . Furthermore, the criteria for absorbing unlearnable modes is not sharp as in our case, but is instead controlled by the ratio of $\kappa$ over the number of datapoints and the eigenfunctions.
4 Renormalization beyond the Gaussian
Here we extend the RG formalism to accommodate non-Gaussian effects associated with the feature modes. Namely, cases where the $\varphi$ random variables (under the data measure) have substantial deviations from normality. We begin with a somewhat artificial but illustrative scenario where the greater features are non-Gaussian yet remain independent from the lesser features. In this case, we show that the previous treatment still holds without any changes. We then proceed with allowing weak non-Gaussianity with potential dependencies between greater and lesser feature modes. In this case, alongside the noise renormalization, we obtain a spatial re-weighting of the mean squared error (MSE) loss or, equivalently, a spatial re-weighting of the ridge parameter.
Considering the first case in which the greater and lesser features are independent, the conditional probability in (22) equals the marginal probability, $P[\varphi_{>}|\varphi_{<}]=P[\varphi_{>}]$, so that the integral over greater modes becomes
$\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}]\exp\!\left\{-\frac{1}{2\sigma^{2}}\left(\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)\right\}\leavevmode\nobreak\ .$ (34)
In accordance with our perturbative RG prescription above, we take $\Phi_{>}$ to contain a single feature mode (the rest are to be integrated out in subsequent steps), which we denote by $\varphi_{q}$, so that $\Phi_{q}$ is the replica space vector whose $m^{\mathrm{th}}$ element is $\Phi_{mq}\coloneqq[\Phi_{q}]_{m}=(f_{mq}-y_{q})\varphi_{q}$, cf. (21). The above integral can then be viewed as the cumulant generating function of $P[\varphi_{q}]\exp\{-\frac{1}{2\sigma^{2}}\Phi_{q}^{\top}\Phi_{q}\!\}$ with $-\Phi_{q}^{\top}\Phi_{<}/\sigma^{2}$ playing the role of the cumulant function argument. Accordingly, the leading order non-Gaussian corrections will enter as $(f_{<}(x)-y_{<}(x))^{4}$ times the fourth cumulant of $\varphi_{q}$. For reasonable (e.g., sub-Gaussian) distributions, [^8] Since our approach is perturbative, we limit our study to distributions without heavy tails, cf. zavatoneveth2021exact . the (2n)th cumulant scales like the variance to the $n^{\mathrm{th}}$ power. The higher cumulants are thus $O\left(\delta c^{2}\right)$ and hence negligible, and we obtain the same equations as in the Gaussian case.
We now turn to the more realistic case in which the lesser and greater feature modes are coupled, and relax the Gaussianity assumption by noting that if the data has a high effective dimension, $d$, then a generic feature – which is a sum of many monomials in $x_{i}$ – can be viewed as a sum of many weakly correlated random variables and hence well-approximated by a Gaussian. This, together with the empirical success of the Gaussian assumption canatar , motivates us to consider corrections to Gaussianity as small $1/d$ corrections (as in tomasini2022failure ) and treat the latter perturbatively. At a technical level, we do so by keeping a leading-order Edgeworth expansion of $P[\varphi_{<},\varphi_{>}]$. We then use this expansion to obtain the conditional distribution $P[\varphi_{>}|\varphi_{<}]$ appearing in (22), and perform the integral perturbatively by keeping only the leading non-Gaussianities. Concretely, we take as our large-$d$ expansion
$P[\varphi_{<},\varphi_{>}]\approx{\cal N}[0,\mathbb{1};\varphi_{<}]{\cal N}[0,\mathbb{1};\varphi_{>}]\left(1+\frac{1}{4!}\sum_{k_{1},k_{2},k_{3},k_{4}}U_{k_{1}k_{2}k_{3}k_{4}}He_{4}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}},\varphi_{k_{4}})\right)$ (35)
where ${\cal N}[0,\mathbb{1};\varphi_{*}]$ denotes a standard normal distribution of the feature mode vector $\varphi_{*}$, $He_{4}$ is the fourth multivariate (probabilist’s) Hermite polynomial,
$He_{4}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}},\varphi_{k_{4}})={\cal N}[0,\mathbb{1};\varphi_{<}]^{-1}{\cal N}[0,\mathbb{1};\varphi_{>}]^{-1}\partial_{\varphi_{k_{1}}}\partial_{\varphi_{k_{2}}}\partial_{\varphi_{k_{3}}}\partial_{\varphi_{k_{4}}}{\cal N}[0,\mathbb{1};\varphi_{<}]{\cal N}[0,\mathbb{1};\varphi_{>}]$ (36)
(see for example Naveh_2021 ), and $U$ is the (fourth) cumulant/Ursell function,
$U_{k_{1}k_{2}k_{3}k_{4}}\coloneqq\langle\varphi_{k_{1}}\varphi_{k_{2}}\varphi_{k_{3}}\varphi_{k_{4}}\rangle_{P[\varphi],\mathrm{connected}}$ (37)
i.e., the connected 4-pt correlation function with respect to $P[\varphi]$. We emphasize that in our perturbative expansion in $1/d$, we shall keep only terms which are leading order in the fourth cumulant.
As per our RG prescription above, we again take $\Phi_{>}$ to consist of a single feature mode $\varphi_{q}$, whose corresponding GP mode $f_{mq}$ has a variance $\delta c$ much smaller than $\sigma_{c}^{2}$. Let us examine the possible fourth cumulants, $U_{k_{1},k_{2},k_{3},k_{4}}$, $U_{k_{1},k_{2},k_{3},q}$, $U_{k_{1},k_{2},q,q}$, $U_{k_{1},q,q,q}$, and $U_{q,q,q,q}$, where henceforth $k_{i}$ refers to lesser modes. Since the variance of the higher mode scales like $\delta c$, the last of these scales like $\delta c^{2}$, while the second to last scales like $\delta c^{3/2}$. Hence, keeping only the leading order terms in $\delta c$ and $1/d$, we arrive at the following RG-equivalent description of (35):
$P[\varphi_{<},\varphi_{q}]$ $\approx{\cal N}[0,\mathbb{1};\varphi_{<}]\left(1+\frac{1}{4!}U_{k1k_{2}k_{3}k_{4}}He_{4}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}},\varphi_{k_{4}})\right)$ (38)
×$\times{\cal N}[0,\mathbb{1};\varphi_{q}]\left(1+\frac{1}{3!}U_{k_{1}k_{2}k_{3}q}He_{3}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}})\varphi_{q}+\frac{1}{2!}U_{k_{1}k_{2}qq}He_{2}(\varphi_{k_{1}},\varphi_{k_{2}})[\varphi_{q}^{2}-1]\right)$
where we have adopted Einstein summation notion, and used relations between multivariate Hermite polynomials implying, e.g., that $He_{4}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}},\varphi_{q})=He_{3}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}})He_{1}(\varphi_{q})$. The combinatorial factors arise from the various ways of fixing $q$. Concretely, the 2nd and 3rd Hermite polynomials for the lesser modes appearing in this expression are
$He_{2}(\varphi_{k_{1}},\varphi_{k_{2}})$ $=\varphi_{k_{1}}\varphi_{k_{2}}-\delta_{k_{1},k_{2}}\leavevmode\nobreak\ ,$ (39)
$He_{3}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}})$ $=\varphi_{k_{1}}\varphi_{k_{2}}\varphi_{k_{3}}-\delta_{k_{1},k_{2}}\varphi_{k_{3}}-\delta_{k_{2},k_{3}}\varphi_{k_{1}}-\delta_{k_{1},k_{3}}\varphi_{k_{2}}\leavevmode\nobreak\ ,$
while
$He_{1}(\varphi_{q})=\varphi_{q}\leavevmode\nobreak\ ,\qquad\qquad He_{2}(\varphi_{q}^{2})=\varphi_{q}^{2}-1\leavevmode\nobreak\ .$ (40)
Finally we note that, since we perform the Edgeworth expansion about the standard normal $\mathcal{N}[0,\mathbb{1};\varphi_{q}]$, both the Hermite polynomials of the higher modes (40) have zero mean under this distribution. Consequently, the top line in (38) is the marginal distribution of the lesser feature modes, and therefore
$P[\varphi_{q}|\varphi_{<}]\equiv\frac{P[\varphi_{q},\varphi_{<}]}{P[\varphi_{<}]}\approx{\cal N}[0,\mathbb{1};\varphi_{q}]\left[1+A\varphi_{q}+B\left(\varphi_{q}^{2}-1\right)\right]\leavevmode\nobreak\ ,$ (41)
where we have introduced the shorthand notation
$A\coloneqq\frac{1}{3!}U_{k_{1}k_{2}k_{3}q}He_{3}(\varphi_{k_{1}},\varphi_{k_{2}},\varphi_{k_{3}})\leavevmode\nobreak\ ,\quad\quad B\coloneqq\frac{1}{2!}U_{k_{1}k_{2}qq}He_{2}(\varphi_{k_{1}},\varphi_{k_{2}})$ (42)
where the first non-trivial term constitutes a shift of the conditional mean, and the second is a shift of the conditional variance of the greater mode.
Now, proceeding from the integral in (22), we will substitute the expansion (41) in place of $P[\varphi_{>}|\varphi_{<}]$. First however, we note that as per our earlier assumption on the relative variance of $\Phi_{>}$, i.e., $\delta c\ll\sigma^{2}$, we may neglect the $\Phi_{>}^{\top}\Phi_{>}$ term in the exponential, so that [^9] Alternatively, one can retain this term throughout the subsequent calculation and show that this amounts to neglecting $O\!\left(\delta c^{2}/\sigma^{4}\right)$ terms.
$\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]\,e^{-\frac{2\Phi_{<}^{\top}\Phi_{>}+\Phi_{>}^{\top}\Phi_{>}}{2\sigma^{2}}}\approx\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]\,e^{-\frac{\Phi_{<}^{\top}\Phi_{>}}{\sigma^{2}}}$ (43)
Additionally, since we take $\Phi_{>}$ to consist of a single feature mode, we keep only a single term in the decomposition (21), so that at each step of the RG, the integral (43) becomes
[…]
Hence, substituting the large-$d$ expansion (41) in place of the conditional distribution and re-exponentiating after performing the Gaussian integral, we obtain
$\int$ $\!\mathcal{D}\varphi_{q}P[\varphi_{q}|\varphi_{<}]e^{-\frac{\Phi_{<}^{\top}\Phi_{q}}{\sigma^{2}}}=\frac{1}{\sqrt{2\pi}}\int\!\mathcal{D}\varphi_{q}e^{-\frac{\varphi^{2}_{q}}{2}-\frac{\Phi_{<}^{\top}\Phi_{q}}{\sigma^{2}}}\left[1+A\varphi_{q}+B\left(\varphi_{q}^{2}-1\right)\right]$ (45)
$=\frac{1}{\sqrt{2\pi}}\int\!\mathcal{D}\varphi_{q}e^{-\frac{\varphi^{2}_{q}}{2}-\frac{1}{\sigma^{2}}\sum_{m}\Phi_{m<}\Phi_{mq}}\left[1+A\varphi_{q}+B\left(\varphi_{q}^{2}-1\right)\right]$
$\approx\exp\left[\frac{1+2B}{2\sigma^{4}}\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}-\frac{A}{\sigma^{2}}\Phi_{<}^{\top}\bm{f}_{q}\right]$
where we have defined $\bm{f}_{q}$ to be the replica-space vector whose $m^{\mathrm{th}}$ element is $[\bm{f}_{q}]_{m}=f_{mq}$, cf. the definition of $\Phi_{q}$ below (34), and the approximation refers to neglecting terms higher-order in $1/d$.
Our next task is to integrate out the greater GP mode ($f_{mq}$), again to leading order in the non-Gaussianity (consistent with our Edgeworth expansion above). Recalling the replicated partition function from Sec. 3, we have
$Z^{M}\right>_{\eta}=e^{-\eta}\int\!\mathcal{D}\bm{f}\,e^{-S_{0}[\bm{f}]+\eta\int\!\mathcal{D}\varphi P[\varphi]\,\exp\left[-\frac{1}{2\sigma^{2}}\left(\Phi_{<}^{\top}\Phi_{<}+2\Phi_{<}^{\top}\Phi_{>}+\Phi_{>}^{\top}\Phi_{>}\right)\right]}$ (46)
where $\mathcal{D}\bm{f}\coloneqq\prod_{m}\mathcal{D}f_{m}=\prod_{mk}\mathrm{d}f_{mk}$ (cf. (16)), and $S_{0}$ denotes the free action given in the first line of (15), which we repeat here for convenience:
$S_{0}[f]=\frac{1}{2}\sum_{m=1}^{M}\sum_{k=1}^{\infty}\frac{1}{\lambda_{k}}|f_{mk}|^{2}\leavevmode\nobreak\ ,$ (47)
By (22) and the above analysis leading to (45), the second term in the exponential of $S_{\mathrm{int}}$ becomes
$S_{\mathrm{int}}$ $=-\eta\int\!\mathcal{D}\varphi P[\varphi]\,\exp\left[-\frac{1}{2\sigma^{2}}\left(\Phi_{<}^{\top}\Phi_{<}+2\Phi_{<}^{\top}\Phi_{>}+\Phi_{>}^{\top}\Phi_{>}\right)\right]$ (48)
$\approx-\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\,\exp\left[-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}+\frac{1+2B}{2\sigma^{4}}\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}-\frac{A}{\sigma^{2}}\,\Phi_{<}^{\top}\bm{f}_{q}\right]\leavevmode\nobreak\ .$
Splitting the path integral as in (16), we thus have
$Z^{M}\right>_{\eta}=e^{-\eta}\int\!\mathcal{D}\bm{f}_{<}\,e^{-S_{0}[\bm{f}_{<}]}\int\!\mathcal{D}\bm{f}_{q}\,e^{-S_{0}[\bm{f}_{q}]+\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\,\exp\left[-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}+\frac{1+2B}{2\sigma^{4}}\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}-\frac{A}{\sigma^{2}}\,\Phi_{<}^{\top}\bm{f}_{q}\right]}\leavevmode\nobreak\ .$ (49)
Now, to compute the path integral over $\bm{f}_{q}$, we again work perturbatively, keeping terms up to order $\bm{f}_{q}^{2}/\sigma^{2}\sim\delta c/\sigma^{2}$, and only linear order in $A$ and $B$ (as per our large-$d$ expansion). As above, this can be done by expanding the interaction terms in the action, then expanding the exponential of the partition function, performing the integral over $\bm{f}_{q}$, and then re-exponentiating the result. To leading order in the non-Gaussianities, we find:
$\int\!\mathcal{D}$ $\bm{f}_{q}\,e^{-S_{0}[\bm{f}_{q}]+\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\,\exp\left[-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}+\frac{1+2B}{2\sigma^{4}}\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}-\frac{A}{\sigma^{2}}\,\Phi_{<}^{\top}\bm{f}_{q}\right]}$ (50)
$\approx$ $\exp\left\{\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\,e^{-\frac{\Phi_{<}^{\top}\Phi_{<}}{2\sigma^{2}}+\lambda_{q}(1+2B)\frac{\Phi_{<}^{\top}\Phi_{<}}{2\sigma^{4}}}\right\}$
where $\lambda_{q}\sim\delta c$, since this is the mode that is integrated-out at a single RG step. As in the Gaussian case, this results in an effective substitution of $\bm{f}_{q}^{\top}\bm{f}_{q}$ and $\bm{f}_{q}$ with their averages under $S_{0}[\bm{f}_{q}]$, namely $\lambda_{q}\mathbb{1}_{M\times M}$ and 00. As expected by the linearity of GP regression, the result is quadratic in $\Phi_{<}$. Interpreting the term in the exponent as a loss function, we see that it remains MSE in nature but becomes “spatially weighted” by the $1+2B[\varphi_{<}]$ factor. Indeed, recasting this term back into $\mathrm{d}\mu_{x}$ and $\phi_{k}(x)$ notation,
$1+2B[\varphi_{<}]$ $\sim 1+2B[\phi_{<}(x)]=1+U_{k_{1}k_{2}qq}He_{2}(\phi_{k_{1}}(x),\phi_{k_{2}}(x))$ (51)
$=1+U_{k_{1}k_{2}qq}\left(\phi_{k_{1}}(x)\phi_{k_{2}}(x)-\delta_{k_{1},k_{2}}\right)\leavevmode\nobreak\ ,$
cf. (20), shows that $\int\!\mathrm{d}\mu_{x}(1+2B(x))=1$ since $B$ is the second cumulant.
Interestingly, this result shows that the spatial weighting of the $L^{2}$ norm undergoes an RG flow. Specifically, after marginalizing over a single mode, we obtain
$\frac{W_{\delta c}(x)}{\sigma^{2}_{\delta c}}=\frac{W_{0}(x)}{\sigma_{0}^{2}}-\frac{\delta c(1+2B_{0}(x))}{\sigma_{0}^{4}}$ (52)
[…]
$\frac{W_{\delta c}(x)}{\sigma^{2}_{0}+\delta c}=\frac{W_{0}(x)}{\sigma_{0}^{2}}-\frac{\delta c[1+2B_{0}(x)]}{\sigma_{0}^{4}}\leavevmode\nobreak\ ,$ (53)
and hence
$W_{\delta c}(x)=W_{0}(x)-\frac{2\,\delta c}{\sigma_{0}^{2}}B_{0}(x)+O\left(\delta c^{2}/\sigma^{4}_{0}\right)\leavevmode\nobreak\ ,$ (54)
where we used the initial condition $W_{0}(x)=1$; notably, $W_{\delta c}(x)$ remains normalized.
Intuitively, the weighting factor is quite natural: it reflects the fact that if, for example, the square of $\phi^{2}_{q}(x)$ has some positive correlations with $\phi^{2}_{k}(x)$ (where $q$ is the higher mode removed at a given RG step, and $k$ is some lesser mode), the marginalization results in varying amounts of effective noise, since more will be generated at spatial locations where large $\phi^{2}_{k}(x)$ is large.
4.1 A solvable toy example: single mode integration
As a concrete application, here we consider perhaps the simplest setting with tractable non-Gaussian effects. However, unlike in the above examples where such effects were treated perturbatively, here we consider a more natural setting relevant to fully connected networks, wherein the lesser feature modes largely determine the greater feature modes. Concretely, we consider $x\in\mathbb{R}$ with $\mathrm{d}\mu_{x}\sim{\cal N}[0,1]$, and a rank-2 kernel of the form
$K(x,y)$ $=\lambda_{1}xy+\lambda_{2}(x^{2}-1)(y^{2}-1)$ (55)
$=\lambda_{1}He_{1}(x)He_{1}(y)+\lambda_{2}He_{2}(x)He_{2}(y)\leavevmode\nobreak\ ,$
where $He_{n}(x)$ is again the probabilist’s Hermite polynomial,
$He_{n}(x)\coloneqq(-1)^{n}e^{\frac{x^{2}}{2}}\frac{\mathrm{d}^{n}}{\mathrm{d}x^{n}}e^{-\frac{x^{2}}{2}}\leavevmode\nobreak\ .$ (56)
in particular, $He_{1}(x)=x$ and $He_{2}(x)=x^{2}-1$. Given the Gaussian measure, $He_{1}(x)$, $He_{2}(x)$ can be shown to span the non-zero eigenspaces of the kernel with eigenvalues ${\lambda_{1}:=\lambda_{<}}$, ${\lambda_{2}:=\lambda_{>}}$ respectively. To connect with our formalism above, note that here we have only two modes, $k\in\{1,2\}$, which are furnished by the first two Hermite polynomials, i.e., $\phi_{1}(x)=He_{1}(x)$ and $\phi_{2}(x)=He_{2}(x)$, cf. (7),
Our task is to learn a target function
$y(x)=x^{5}-10x^{3}+15x=He_{5}(x)\leavevmode\nobreak\ .$ (57)
From a spectral bias perspective canatar , the average predictor for this model is zero, since the target is orthogonal to all non-zero eigenspaces under the input measure. [^11] By this we mean that the inner product on the associated Hilbert space vanishes, $\int\!\mathrm{d}\mu_{x}\,K(x,y)He_{5}(x)=0$. This is not necessarily the case however, since the features ($\varphi$, defined in (20)) will be non-Gaussian unless $\phi(x)$ is an affine function of $x\sim\mathcal{N}[0,1]$. Statistically therefore, the more common events in which $x$ lies closer to zero may bias the predictor on small datasets to act as if it were seeing the function $-10x^{3}+15x$, which does have non-vanishing overlap with $He_{1}(x)$. We show below that our RG approach can accurately predict such effects, and that they can be viewed as stemming from an MSE loss weighted by $\left(1-\lambda_{2}\sigma^{-2}He_{2}(x)^{2}\right)$. Indeed, while $y=He_{5}(x)$ and $He_{1}(x)$ are not coupled by standard MSE loss, the extra weighting term yields a non-zero overlap via integrals of the type $\int\!\mathrm{d}\mu_{x}He_{1}(x)\left(\lambda_{2}\sigma^{-2}He_{2}(x)^{2}\right)He_{5}(x)$.
We shall first derive the above weighted form of the MSE loss, and then discuss its effect on predictions. Since the kernel mode associated with $He_{5}(x)$ is frozen/non-fluctuating, we do not treat it as a GP mode, and instead integrate out only the mode associated with $\lambda_{2}$. [^12] That is, we view $\bm{y}_{<}(x)$ as $\bm{y}_{1}He_{1}(x)+\bm{y}_{5}He_{5}(x)=He_{5}(x)$, with $y_{1}=0$ and $y_{5}=1$, and similarly $f_{<}(x)=f_{1}x$ with $f_{5}=0$. Alternatively, one may treat $y$ as part of the greater target modes, where its effect would then enter as a subleading shift to the lesser target. We denote this by $\varphi_{>}=\varphi_{2}$, and the associated replica-space GP mode by $\bm{f}_{2}$, with $\bm{y}_{2}=0$ (i.e., $q=2$ in our notation above). Furthermore, we take ${\bm{f}}_{1}$ to be the replica space vector associated with the linear (lesser) mode and (i.e., $He_{1}(x)$) and ${\bm{y}}_{1}$ to be the corresponding target. Converting to feature variables $\varphi_{<}=\varphi_{1}$ and $\varphi_{>}=\varphi_{2}$ via (20), we have
$P[\varphi_{1}]$ $=\int\mathrm{d}\mu_{x}\delta\left(\varphi_{1}-x\right)=\mathcal{N}[0,\mathbb{1};\varphi_{1}]\leavevmode\nobreak\ ,$ (58)
$P[\varphi_{1},\varphi_{2}]$ $=P[\varphi_{1}]P[\varphi_{2}|\varphi_{1}]=\mathcal{N}[0,\mathbb{1};\varphi_{1}]\int\!\mathrm{d}\mu_{x}\delta\left(\varphi_{2}-He_{2}(x)\right)$
$=\mathcal{N}[0,\mathbb{1};\varphi_{1}]\delta\left(\varphi_{2}-He_{2}(\varphi_{1})\right)=\mathcal{N}[0,\mathbb{1};\varphi_{1}]\delta\left[\varphi_{2}-\left(\varphi_{1}^{2}-1\right)\right]\leavevmode\nobreak\ .$
Notably, unlike the previously discussed perturbative case where $\phi_{>}$ was weakly affected by $\phi_{<}$, here $\phi_{<}$ fully determines $\phi_{>}$. (This scenario is relevant for real-world networks, since the same deterministic behavior is expected for any fully connected network with $\mathrm{d}\mu_{x}$ as the uniform measure on the hypersphere, where the lesser linear modes would determine all greater modes.)
Now, consider the interaction term in the replicated partition function from (22). As a result of the delta function, the integral over the greater feature mode can be easily resolved to yield
$S_{\mathrm{int}}$ $=-\eta\int\!\mathrm{d}\varphi_{1}\,P[\varphi_{1}]e^{-\frac{1}{2\sigma^{2}}\Phi_{1}^{\top}\Phi_{1}}\int\!\mathrm{d}\varphi_{2}P[\varphi_{2}|\varphi_{1}]\exp\!\left\{-\frac{1}{2\sigma^{2}}\left(\Phi_{2}^{\top}\Phi_{2}+2\Phi_{1}^{\top}\Phi_{2}\right)\right\}$ (59)
$=-\eta\int\!\mathrm{d}\varphi_{1}\,P[\varphi_{1}]e^{-\frac{1}{2\sigma^{2}}\Phi_{1}^{\top}\Phi_{1}}\int\!\mathrm{d}\varphi_{2}\delta\left(\varphi_{2}-(\varphi_{1}^{2}-1)\right)e^{-\frac{1}{2\sigma^{2}}\left(\varphi_{2}^{2}|\bm{f}_{2}|^{2}+2\varphi_{2}\varphi_{1}\bm{f}_{2}^{\top}(\bm{f}_{1}-\bm{y}_{1})\right)}$
$=-\eta\int\!\mathrm{d}\varphi_{1}\,P[\varphi_{1}]\exp\left\{-\frac{1}{2\sigma^{2}}\left(\Phi_{1}^{\top}\Phi_{1}+He_{2}(\varphi_{1})^{2}|\bm{f}_{2}|^{2}+2He_{2}(\varphi_{1})He_{1}(\varphi_{1})\bm{f}_{2}^{\top}(\bm{f}_{1}-\bm{y}_{1})\right)\right\}$
Next, we must integrate the greater GP mode $f_{2}$ in the replicated partition function (46), $Z^{M}\right>_{\eta}=e^{-\eta}\int\!\mathcal{D}\bm{f}\,e^{-S_{0}[\bm{f}]-S_{\mathrm{int}}}$. As in (50), we capitalize on the fact that $f_{2}$ scales as $\sqrt{\lambda_{2}}$, and keep all contributions below order $f_{2}^{4}=O(\lambda_{2}^{2})$. After expanding both exponentials, performing the integral, and re-exponentiating, we again find that the effect is to replace $f_{2}$ by the variance under $S_{0}[\bm{f}_{2}]$, [^13] Here we have also dropped the term $M\frac{\lambda_{q}}{2\sigma^{2}}He_{2}(\varphi_{1})^{2}$ that arises from $|\bm{f}_{q}|^{2}=\mathrm{tr}(\bm{f}_{q}^{\top}\bm{f}_{q})\rightarrow M\lambda_{q}$, since this vanishes in the replica limit. Such a factor of $M$ did not arise in (50), since there $\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}\sim\bm{f}_{q}\bm{f}_{q}^{\top}$ does not involve a trace over replica indices.
$Z^{M}\right>_{\eta}$ $=\int\!\mathcal{D}f_{1}\exp\left\{-S_{0}[\bm{f}_{1}]+\eta\int\!\mathrm{d}\varphi_{1}P[\varphi_{1}]e^{-\frac{1}{2\sigma^{2}}\Phi_{1}^{\top}\Phi_{1}\left(1-\frac{\lambda_{2}}{\sigma^{2}}He_{2}(\varphi_{1})^{2}\right)}\right\}$ (60)
$=\int\!\mathcal{D}f_{1}\exp\left\{-S_{0}[\bm{f}_{1}]+\eta\int\!\mathrm{d}\mu_{x}\,e^{-\frac{1}{2\sigma^{2}}\Phi_{1}^{\top}\Phi_{1}\left(1-\frac{\lambda_{2}}{\sigma^{2}}He_{2}(x)^{2}\right)}\right\}\leavevmode\nobreak\ ,$
where on the second line we used (58) to replaced $\varphi_{1}$, $\int\!\mathrm{d}\varphi_{1}P[\varphi_{1}]$ with $x$, $\int\!\mathrm{d}\mu_{x}$, respectively. We thus observe the advertised reweighting of the MSE term, namely
[…]
Next, we perform an EK-type approximation on (60). Namely, we assume $\sigma^{2}$ is large enough compared to $\left(\bm{f}_{1}(x)-\bm{y}_{1}(x)\right)$ to perform a Taylor expansion of the exponent appearing in the action. We note in passing that the next-to-leading order in this expansion Cohen_2021 . happens to vanish for $He_{n\geq 5}(x)$ (but not, e.g., for $He_{3}(x)$). Within the leading order expansion, the action in (60) becomes
$\frac{1}{2\lambda_{1}}\bm{f}_{1}^{2}-\eta\int\mathrm{d}\mu_{x}\left(1-\frac{1}{2\sigma^{2}}(\bm{f}_{1}x-He_{5}(x))^{2}\left[1-\frac{\lambda_{2}}{\sigma^{2}}He_{2}(x)^{2}\right]\right)\leavevmode\nobreak\ .$ (62)
Without the weighting factor, the extremum of $f_{1}$ would be zero here as $He_{1}(x)=x$ and $He_{5}(x)$ do not overlap under $\int\mathrm{d}\mu_{x}$ (since only the cross-term, which is linear in $\bm{f}_{1}$, affects the average). However, with the weighting factor, and using $\int\mathrm{d}\mu_{x}\,xHe_{5}(x)He_{2}(x)^{2}=120$, the saddlepoint is now
$\bar{f}_{1}\approx-\frac{\lambda_{1}}{\lambda_{1}+\sigma^{2}/\eta}\frac{120\lambda_{2}}{\sigma^{2}}\leavevmode\nobreak\ .$ (63)
where the approximation denotes leading order in $\lambda_{2}\sim\delta c$.
To test this empirically, we sampled various draws of datasets with $n=100$ from $x\sim{\cal N}[0,1]$, performed GP regression on these with the above kernel using a fixed random test set with $n_{\mathrm{test}}=1000$, and averaged the results over the various draws. Doing so for $\lambda_{2}=0$ shows an almost zero outcome, as expected. However, taking finite $\lambda_{2}$ reveals a trend that spectral bias Cohen_2021 ; canatar misses. Notably, the fact that our toy example involves only one mode which is integrated out, together with the need to take its variance to be small ($\lambda_{2}/\sigma^{2}\ll 1$) as well as make it unlearnable ($n\lambda_{2}/\sigma^{2}\ll 1$) results in a rather small effect. Still, it serves to demonstrate the conceptual phenomena whose importance depends on the setting or the desired accuracy.
Figure 1: Non-Gaussian features and spatial re-weighting effects. Theory versus experiment for the model of sec. 4.1 with $n=100$ datapoints $\sigma^{2}=400$ and $\lambda_{>}:=\lambda_{2}=0.1$ (unless stated otherwise). Learning a 5th Hermite polynomial, using a kernel capable of expressing only 1st and 2nd Hermite polynomials should give a zero average predictor (green line) based on the standard theory canatar ; Cohen_2021 . However, due to spatial re-weighing, a coupling between 1st and 5th Hermite polynomial arises leading to a non-zero result.
5 Discussion
In this work, we applied the powerful formalism of Wilsonian RG to Gaussian Process regression. In particular, treating the output of a trained DNN as a statistical ensemble, we split the associated fluctuations into high- and low-energy modes (based on their stiffness in a random network). Provided the variance of the higher modes is much smaller than the ridge parameter, these may be integrated out in a tractable manner leading to a renormalized GP regression problem. In simple settings where the features associated with those modes are Gaussian, this results in a renormalized ridge parameter. In more general settings, one obtains an effective GP regression problem where the ridge parameter becomes spatially dependent. We then applied this to a simple toy model, in which the effects determined via our RG approach are demonstrably missed in the standard eigen-learning framework canatar .
It would be interesting to generalize our formalism to include fluctuation effects over datasets by keeping track of $y_{>}^{2}$ dependencies. Furthermore, it would be interesting to extend our one-step RG treatment for the non-Gaussian case into a fully-fledged functional RG flow. This may enable us to define a notion of relevancy and irrelevancy applicable to generic non-Gaussian features and, potentially, a notion of universality classes for DNNs. Similarly, our formalism could be extended to finite neural networks where additional interaction terms would appear corresponding to finite-width corrections to the Gaussian prior, and investigate universality in those corrections.
Acknowledgements
JNH was supported by the National Science Foundation under Grant No. NSF PHY-1748958 and by the Gordon and Betty Moore Foundation through Grant No. GBMF7392. AM acknowledges support from Perimeter Institute, which is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development and by the Province of Ontario through the Ministry of Colleges and Universities. ZR acknowledges support from ISF Grant 2250/19. JNH, RJ, and AM wish to thank the Mainz Institute for Theoretical Physics for hosting the 2022 workshop “A Deep-Learning Era of Particle Theory”, where this collaboration was initiated, as well as the Aspen Center for Physics (ACP, supported by National Science Foundation grant PHY-2210452) for hospitality during the 2023 winter conference “Theoretical Physics for Machine Learning”. JNH, and ZR acknowledge further support from ACP through the 2023 summer program “Theoretical Physics for Deep Learning”. This research was also supported in part by grants NSF PHY-1748958 and PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP), through the participation of AM, JNH, and ZR in the Fall 2023 program “Deep Learning from the Perspective of Physics and Neuroscience.”
Appendix A Spectral decomposition of the Gaussian action
In this appendix, we perform the spectral decomposition given in (9) on the quadratic action (6), repeated here for convenience:
$S=\sum_{m=1}^{M}\frac{1}{2}\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\,f_{m}(x)K^{-1}(x,x^{\prime})f_{m}(x^{\prime})+\frac{\eta}{\sigma^{2}}\int\!\mathrm{d}\mu_{x}\sum_{m}\frac{\left(f_{m}(x)-y(x)\right)^{2}}{2}\leavevmode\nobreak\ .$ (64)
Beginning with the first term, we have
$\int$ $\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\,f_{m}(x)K^{-1}(x,x^{\prime})f_{m}(x^{\prime})$ (65)
$=\int\!\mathrm{d}\mu_{x}\mathrm{d}\mu_{x^{\prime}}\left(\sum_{k}f_{mk}\phi_{k}(x)\right)K^{-1}(x,x^{\prime})\left(\sum_{k^{\prime}}f_{mk^{\prime}}\phi_{k^{\prime}}(x^{\prime})\right)$
$=\sum_{k,k^{\prime}}f_{mk}f_{mk^{\prime}}\!\int\!\mathrm{d}\mu_{x}\phi_{k}(x)\left(\int\!\mathrm{d}\mu_{x^{\prime}}K^{-1}(x,x^{\prime})\phi_{k^{\prime}}(x^{\prime})\right)$
$=\sum_{k,k^{\prime}}\lambda_{k^{\prime}}^{-1}f_{mk}f_{mk^{\prime}}\!\int\!\mathrm{d}\mu_{x}\,\phi_{k}(x)\phi_{k^{\prime}}(x)$
$=\sum_{k,k^{\prime}}\lambda_{k^{\prime}}^{-1}f_{mk}f_{mk^{\prime}}\delta_{kk^{\prime}}=\sum_{k}\frac{1}{\lambda_{k}}|f_{mk}|^{2}\leavevmode\nobreak\ ,$
where in going to the fourth line we have used (7), and the last line follows from the orthonormality relation (8). Meanwhile, in the second term we have
$\int\!\mathrm{d}\mu_{x}\left(f_{m}(x)-y(x)\right)^{2}$ $=\int\!\mathrm{d}\mu_{x}\left[\sum_{k}\left(f_{mk}-y_{k}\right)\phi_{k}(x)\right]\left[\sum_{k^{\prime}}\left(f_{mk^{\prime}}-y_{k^{\prime}}\right)\phi_{k^{\prime}}(x)\right]$ (66)
$=\sum_{k,k^{\prime}}\left(f_{mk}-y_{k}\right)\left(f_{mk^{\prime}}-y_{k^{\prime}}\right)\!\int\!\mathrm{d}\mu_{x}\,\phi_{k}(x)\phi_{k^{\prime}}(x)$
$=\sum_{k}\left(f_{mk}-y_{k}\right)^{2}\leavevmode\nobreak\ ,$
where the last step again follows from (8). Substituting these into (64) then yields the decoupled feature-space action (10), as desired.
Note that these expressions will be unchanged when performing the split into greater and lesser GP modes in (13), since the cross-terms will vanish by the orthonormality relation (8). For example, the fourth line of (65) becomes
$\int$ $\mathrm{d}\mu_{x}\left(\sum_{k\leq\kappa}f_{mk}\phi_{k}+\sum_{k>\kappa}f_{mk}\phi_{k}\right)\left(\sum_{k^{\prime}\leq\kappa}\lambda_{k^{\prime}}^{-1}f_{mk^{\prime}}\phi_{k^{\prime}}+\sum_{k^{\prime}>\kappa}\lambda_{k^{\prime}}^{-1}f_{mk^{\prime}}\phi_{k}^{\prime}\right)$ (67)
$=\left(\sum_{k,k^{\prime}\leq\kappa}+\sum_{k,k^{\prime}>\kappa}+\sum_{\begin{subarray}{c}k\leq\kappa\\ k^{\prime}>\kappa\end{subarray}}+\sum_{\begin{subarray}{c}k>\kappa\\ k^{\prime}\leq\kappa\end{subarray}}\right)\lambda_{k^{\prime}}^{-1}f_{mk}f_{mk^{\prime}}\int\!\mathrm{d}\mu_{x}\,\phi_{k}(x)\phi_{k^{\prime}}(x)$
$=\sum_{k}\frac{1}{\lambda_{k}}|f_{mk}|^{2}$
where the last two summations vanish since $k\!\neq k^{\prime}$.
Appendix B Simplifying interactions in the replica limit
The replicated partition function allows us to calculate observables by adding some source term (e.g., $\sum_{k}\alpha_{k}f_{k}$) and taking functional derivatives,
$\partial_{\alpha}\langle\ln(Z_{\alpha})\rangle_{\eta}\mathrel{\mathop{=}\limits_{M\rightarrow 0}}\partial_{\alpha}\frac{\langle Z_{\alpha}^{M}\rangle_{\eta}-1}{M}\leavevmode\nobreak\ .$ (68)
Such a procedure has a predictable dependence on $y$. In particular, since each $\ln(Z_{\alpha})$ describes the Gaussian Process/distribution for a given training set, only $\partial_{\alpha}\ln(Z_{\alpha})$ scales as $y$, while higher derivatives/cumulants do not. This scaling survives dataset averaging. Consequently, we may imagine an infinitesimal scaling variable $\epsilon$, take $y\rightarrow\epsilon y$, and remember to divide by $\epsilon^{-1}$ the average predictor (but not any of the higher cumulants). Equivalently, we may ignore any $y^{2}$ or higher contribution, and use the following equivalent version of $C$:
$C_{mm^{\prime}}\rightarrow\sum_{k>\kappa}\left(f_{km}f_{km^{\prime}}+y_{k}(f_{km}+f_{km^{\prime}})\right)$ (69)
cf. (26).
Our next aim is to show that perturbation theory in $C$, where $C$ appears to the $n^{\mathrm{th}}$ order, yields terms which scale as $\big{(}\delta c/\sigma_{c}^{2}\big{)}^{n}$. The power of $\sigma_{c}^{2}$ stems from the simple fact that $C$ always appears accompanied by $\sigma^{2}$ in the interaction, which becomes the effective parameter $\sigma_{c}^{2}$ under RG. Meanwhile, the power of $\delta c$ can be understood from the fact that, had we not included the $y_{k}$ component in $C_{mm^{\prime}}$, the two factors of $f_{mk}$ from perturbation theory with respect to the free theory for the greater modes (cf. $S_{0}[f_{>}]$ in (10)) would contract with the same GP mode, thereby yielding the desired power of $\delta c$. Turning to the $y_{k}$ term may then be treated as adding a $\sqrt{\delta c}$ factor, since the single factor of $f_{mk}$ appearing in (69) averages to zero and must get contracted. However, due to our argument about $y$ above, it may only appear to first power overall in the perturbative expansion (indeed higher power would yield a free replica summation which would nullify in the replica limit). Furthermore, being antisymmetric in the greater modes ($f_{mk}$), it may only appear when calculating observables with an odd power of greater modes. Subsequently, if we are interested only in lesser observables we may take $y_{>}=0$ in the derivation for Gaussian features.
Using the above reasoning and our previous infinitesimal (i.e., leading order $\delta c/\sigma^{2}$) approach, we may integrate out the greater GP modes using leading order perturbation theory in $C$. As mentioned in the main text, this essentially amounts to replacing $C$ by its average under the free theory wherever it appears in the interaction. As this yields a $\delta c\mathbb{1}_{M\times M}$ with no $y_{>}$ dependence, we may take $y_{>}=0$ from the start in the Gaussian case when calculating lesser observables. The effect of this procedure on the “loss” part of the interaction is discussed in the main text. This procedure also exposes that the prefactor $\exp\left\{-\frac{1}{2}\mathrm{tr}\ln\left(C/\sigma^{2}+\mathbb{1}\right)\right\}$ appearing in (28) becomes $\exp\left\{-\frac{1}{2}\mathrm{tr}\ln\left(\left(\delta c/\sigma^{2}+1\right)\mathbb{1}_{M\times M}\right)\right\}=\exp\left\{-\frac{M}{2}\mathrm{tr}\ln\left(\delta c/\sigma^{2}+1\right)\right\}$, which goes to 1 in the replica ($M\to 0$) limit. This justifies its removal in the main text.
The same general reasoning applies to the non-Gaussian case, however extra care is needed for the terms proportional to $A$ as these contain a single power of $y$ which may affect the average predictor.